A vendor sells hot dogs, bags of potato chips, and soft drinks. A customer buys 3 hot dogs, 4 bags of potato chips, and 5 soft drinks for $12.50. The price of a hot dog is $0.75 more than the price of a bag of potato chips. The cost of a soft drink is $2.00 less than the price of two hot dogs. Find the cost of each item.
Answer by stanbon(75887)

A vendor sells hot dogs, bags of potato chips, and soft drinks. A customer buys 3 hot dogs, 4 bags of potato chips, and 5 soft drinks for $12.50. The price of a hot dog is $0.75 more than the price of a bag of potato chips. The cost of a soft drink is $2.00 less than the price of two hot dogs. Find the cost of each item.
-------------
3h + 4c + 5d = 12.50
h = c + 0.75
d = 2h-2
---------
Substitute for "c" and "d" and solve for "h"::
3h + 4(h-0.75) + 5(2h-2) = 12.50
17h - 7 = 12.50
17h = 19.50
h = $1.15 (price of a hot dog)
-----
c = h-.75 = 0.40 (price of bag of chips)
-----
d = 2h-2 = 0.30 (price of a soft drink)
-----------
Cheers,
Stan H.
---------------